# Sliding Ferroelectrics Induced Hybrid-Order Topological Phase Transitions

### v2

@@ Abstract
Approach &\&& Model— We consider a spin- and layer-locked bilayer AFM second order topological insulator (SOTI). The interlayer sliding simultaneously generates ferroelectric polarization and breaks the $\{C_{2}||m_{z}\}$ symmetry, which consequently induces spin splitting, as illustrated in Fig. 1 (a). Fig. 1 (b) shows that SHTI arises from spin-selective band inversion driven by polarization potential during sliding. The process starts from a SOTI with nonzero corner charge ($Q_{s}$) but vanishing spin Chern number ($C_{s}$). Upon increasing polarization potential, a band inversion occurs in one spin channel, rendering both $Q_{s}$ and $C_{s}$ nonzero. The sliding ferroelectric polarization is also limited, as can be found in the general theory proposed by Ji et al. [55]. Among the many 2D higher-order topological materials, systems with hexagonal lattices have a distinct advantage.
@@
where $Q^{\uparrow/\downarrow}$ and $C^{\uparrow/\downarrow}$ represent the spin corner charge and the spin Chern number, respectively. The system’s second-order topological index $Q_{s}$ is protected by the $C_{3}$ rotational symmetry [58]. For the parity calculation of all occupied states at high-symmetry points in the Brillouin zone, one can take $[\mathrm{K}_{n}^{(3)}]=\#\mathrm{K}_{n}^{(3)}-\#\Gamma_{n}^{(3)}$, where ##\## denotes counting with respect to the symmetry eigenvalues at $K$ and $\Gamma$ points. The eigenvalues of the $C_{3}$ rotation are defined as $e^{2\pi i(n-1)/3}\left(\text{for}\ n=1,2,3\right)$. The topological index of higher-order topological crystalline insulators is $\chi^{(3)}=([K_{1}^{(3)}],[K_{2}^{(3)}]),\ Q_{s}^{(3)}=\frac{e}{3}[K_{2}^{(3)}]\text{mod}\ e$, where $e$ represents the charge of a free electron [58, 19]. For spin-distinguished magnetic materials, spin corner charge $Q_{s}^{(3)}=e/3$.
@@
To compare different topological phases, Fig. 3 shows band projection diagrams for four phases. Initially, the valence band at the $K$ valley is mainly from $d_{xy}$ and $d_{x^{2}-y^{2}}$ orbitals, characteristic of SOTIs before the topological phase transition [18]. In the triangular quantum dot under open boundary conditions, in-gap eigenvalues correspond to corner-localized states on both layers, as in Fig. 3 (a). Band inversion first emerges at the $K$ valley for spin-up states as the nearest-neighbor hopping decreases, with the valence band maximum switching from $d_{xy}/d_{x^{2}-y^{2}}$ to $d_{z^{2}}$ orbitals, as shown in Fig. 3(f), rendering spin-up into a Chern topological insulator while spin-down remains a SOTI. The band-gap eigen wave functions reveal a SHTI, as shown in Fig. 3 (b). Subsequently, spin-down undergoes a band inversion at the $K^{\prime}$ valley. With Chern numbers of $+1$ and $-1$ for the upper and lower layers, respectively, the system enters the QSHI, as shown in Fig. 3 (c). Another band inversion then occurs in the upper layer with $C_{s}=0$, driving the system into the QAHI, as shown in Fig. 3 (d). Finally, as spin-down transitions into a NI, the system reaches a trivial phase.

### v3

@@ Abstract
Approach &\& Model— We consider a spin- and layer-locked bilayer AFM second order topological insulator (SOTI). The interlayer sliding simultaneously generates ferroelectric polarization and breaks the $\{C_{2}||m_{z}\}$ symmetry, which consequently induces spin splitting, as illustrated in Fig. 1 (a). Fig. 1 (b) shows that SHTI arises from spin-selective band inversion driven by polarization potential during sliding. The process starts from a SOTI with nonzero corner charge ($Q_{s}$) but vanishing spin Chern number ($C_{s}$). Upon increasing polarization potential, a band inversion occurs in one spin channel, rendering both $Q_{s}$ and $C_{s}$ nonzero. The sliding ferroelectric polarization is also limited, as can be found in the general theory proposed by Ji et al. [55]. Among the many 2D higher-order topological materials, systems with hexagonal lattices have a distinct advantage.
@@
where $Q^{\uparrow/\downarrow}$ and $C^{\uparrow/\downarrow}$ represent the spin corner charge and the spin Chern number, respectively. The system’s second-order topological index $Q_{s}$ is protected by the $C_{3}$ rotational symmetry [58]. For the parity calculation of all occupied states at high-symmetry points in the Brillouin zone, one can take $\mathrm{K}_{n}^{(3)}]=\#\mathrm{K}_{n}^{(3)}-\#\Gamma_{n}^{(3)}$, where #\# denotes counting with respect to the symmetry eigenvalues at $K$ and $\Gamma$ points. The eigenvalues of the $C_{3}$ rotation are defined as $e^{2\pi i(n-1)/3}\left(\text{for}\ n=1,2,3\right)$. The topological index of higher-order topological crystalline insulators is $\chi^{(3)}=([K_{1}^{(3)}],[K_{2}^{(3)}]),\ Q_{s}^{(3)}=\frac{e}{3}[K_{2}^{(3)}]\text{mod}\ e$, where $e$ represents the charge of a free electron [58, 19]. For spin-distinguished magnetic materials, spin corner charge $Q_{s}^{(3)}=e/3$.
@@
To compare different topological phases, Fig. 3 shows band projection diagrams for four phases. Initially, the valence band at the $K$ valley is mainly from $d_{xy}$ and $d_{x^{2}-y^{2}}$ orbitals, characteristic of SOTIs before the topological phase transition [18]. In the triangular quantum dot under open boundary conditions, in-gap eigenvalues correspond to corner-localized states on both layers, as in Fig. 3 (a). Band inversion first emerges at the $K$ valley for spin-up states as the nearest-neighbor hopping decreases, with the valence band maximum switching from $d_{xy}/d_{x^{2}-y^{2}}$ to $d_{z^{2}}$ orbitals, as shown in Fig. 3(f), rendering spin-up into a Chern topological insulator while spin-down remains a SOTI. The band-gap eigen wave functions reveal a SHTI, as shown in Fig. 3 (b). Subsequently, spin-down undergoes a band inversion at the $K^{\prime}$ valley. With Chern numbers of +1+1 and 1-1 for the upper and lower layers, respectively, the system enters the QSHI, as shown in Fig. 3 (c). Another band inversion then occurs in the upper layer with $C_{s}=0$, driving the system into the QAHI, as shown in Fig. 3 (d). Finally, as spin-down transitions into a NI, the system reaches a trivial phase.